if the total height of a building us h and a ball covers a distance of 9h/25 during the last second of its journey when it ie total heingts drpped from the top, calculate the total height

Grade:11

1 Answers

SOURAV MISHRA
10 years ago
g(2t - 1)/2 = 9h/25gt^2/2 = hthese equations give(2t - 1)/t^2 = 9/25this implies that 9t^2 - 50t + 25 = 0the roots are t = 5 and 5/9.t = 5/9 is not acceptable as it is not an integer so the time taken is 5s.therefore the height of the building is 25g/2 = 125m. (taken g = 10m/s^2)